# Meta-Generalized-Gradient Approximation made Magnetic

### v1

@@ Abstract
The Jacob’s ladder of density functional theory (DFT) proposes the compelling view that by extending the form of successful approximations — being guided by exact conditions and selected (least empirical) norms — upper rungs will do better than the lower, thus allowing to balance accuracy and computational effort. Meta-generalized-gradient-approximations (MGGA) belong to the last rung of the semi-local approximations before hybridization with non-local wave function theories. The performance of MGGAs on non magnetic systems greatly improves upon that of GGAs from the lower rung. Instead, failures of MGGAs on magnetic materials are daunting and make GGAs more reliable for magnetism. This raises the question as to whether consistent improvements can be achieved for magnetism along the lines of the Jacob’s ladder. Here, we present a solution from first principles: departing from the standard approach (i.e. collinear spin-DFT), we exploit the U(1)$\times$×SU(2) gauge symmetry of the exact functional of spin-current-DFT to access an extended set of building blocks for approximations. Crucially, this allows us to treat broken symmetry ferromagnetic and antiferromagnetic states. For these states, we derive a collinear MGGA that performs well for prototypical magnets and make it available in an implementation in the Crystal electronic structure package.
@@
Introduction.  Despite nearly one hundred years of quantum physics, magnetism keeps challenging our understanding of electron correlations. An unsettling example is given by the recently developed strongly-constrained and appropriately normed (SCAN) meta-generalized gradient approximation for density functional theory (DFT).Sun et al. (2015, 2016)
@@
SCAN satisfies a large set of exact constraints, from known properties of the exact exchange-correlation (xc) functional. As a consequence of which, the SCAN gives accurate results for many materials at a computational cost which is comparable with the long known generalized-gradient-approximation (GGA). Therefore it can be considered a modern milestone from which it is possible to develop more accurate approximations.Aschebrock and Kümmel (2019)
@@
But SCAN does not reproduce correctly the stability and properties of prototypical ferromagnets like Fe.Fu and Singh (2018); Ekholm et al. (2018); Fu and Singh (2019); Mejía-Rodríguez and Trickey (2019); Tran et al. (2020) Overestimated tendency toward magnetism and exaggeration of magnetic energies are obtained for transition metal systems. Extension of the analysis to antiferromagnetic solids has shown that modifications of meta-GGAs (MGGAs) yielding an improved description of ferromagnetic metals produce a deterioration in antiferromagnetic insulators, and vice versa.Tran et al. (2020) The status is such that MGGAs yield no advantage over lower rungs of the semi-local approximation.Romero and Verstraete (2018); Tran et al. (2020)
@@
Here, we argue that the problem has persisted due to the fact that the most important exact constraint for dealing with magnetism has not guided most of the DFT developments thus far. We resolve this issue by switching from collinear spin-DFTvon Barth and Hedin (1972); Gunnarsson and Lundqvist (1976) (the current standard methodology) to the fully non-collinear Vignale-Rasolt-Bencheikh spin-current DFT (SCDFT) – a complete DFT theory for dealing with magnetism within the framework of the two-component Pauli-like equation.Vignale and Rasolt (1988); Bencheikh (2003) Remarkably, this extended framework can also solve problems for collinear DFT.
@@
In SCDFT, the most distinctive property of the exact exchange-correlation (xc) energy functional is the form invariance under U(1)$\times$×SU(2) gauge transformations.Vignale and Rasolt (1988); Bencheikh (2003); Pittalis et al. (2017) These transformations rotate spinors as well as the phase of each component at each point in space [see Eq. (11) below]. Failing to satisfy this invariance leads to functional forms that do not necessarily distinguish between physical variations of the spin density and those that arise from a mere change of reference frame. Thus, we argue that the gauge invariance of the exact xc-energy functional must be included as a most important exact constraint in guiding the construction of approximate functionals.
The need to work within the larger scheme of SCDFT arises from the ubiquitous phenomenon of spontaneous symmetry breaking. All magnetic states of matter spontaneously break SU(2) symmetry. Working within the general SCDFT, such low symmetry solutions can be described while gaining extra building blocks for the construction of extended approximate functional forms. As shown below, this feature of the formalism allows us to derive a modified SCAN that performs well also for prototypical magnets.
@@
Recap of the State-of-the-Art.  DFT can in principle find the particle density and total energy of real interacting ground states by solving (self-consistently) an auxiliary non-interacting problem
@@
which is the celebrated Kohn-Sham equation. This is a Schrödinger-like equation that describes single-orbitals corresponding to the actual external potential (the second term on the l.h.s.) added to the Hartree potential (the third term on the l.h.s.) and the xc potential, $v_{\rm xc}$. The latter potential is such that the actual particle density is retrieved $n=\sum^{occ}_{i}|\varphi_{i}|^{2}$not (a) Correspondingly, the energy of the ground state of the real system $E$ is thus obtained by
@@
where $T_{s}=\frac{1}{2}\sum_{i}^{\rm occ}\int d^{3}r~{}|\nabla\varphi_{i}|^{2}$ is the Kohn-Sham kinetic energy, $E_{\rm H}[n]=\int\int d^{3}rd^{3}r^{\prime}\frac{n({\bf r})n({\bf r}^{\prime})}{2|{\bf r}-{\bf r}^{\prime}|}$ is the Hartree energy and $E_{\rm xc}[n]$ is the xc energy functional. For retrieving the exact total energy of the actual state, $E_{\rm xc}[n]$ accounts for the electron-electron energy beyond the classical evaluation obtained via $E_{\rm H}[n]$ as well as the kinetic energies beyond $T_{s}$. Also notice that $v_{\rm xc}=\delta E_{\rm xc}/\delta n$. All that is needed to make the above approach practical is a (numerically amenable) density functional approximation (DFA) for $E_{\rm xc}[n]$. Solutions to $n$ and $E$ can thus be found in an iterative fashion.
The first and simplest $E^{\rm DFA}_{\rm xc}[n]$ is the local density approximation (LDA): Kohn and Sham (1965)
@@
where the xc energy density is approximated locally with that of the uniform gas, $\epsilon^{\rm unif}_{\rm xc}(n)$, for which analytical and accurate numerical reference data are available. To capture the variations of more realistic states, a generalized-gradient approximation (GGA) uses information retrieved from the gradients of the density via $s=|\nabla n|/\left[2(3\pi^{2})^{1/3}n^{4/3}\right]$:
@@
where $F_{\rm xc}^{\rm GGA}(n,s)$ is the xc enhancement factor that connects the LDA energy densities to the GGA energy densities. For slowly varying density, the LDA is retrieved (i.e. $F_{\rm xc}^{\rm GGA}(n,0)=1$). The most popular choice for the enhancement factor is the one by Perdew, Burke and Ernzerhof $F_{\rm xc}^{\rm PBE}$. Perdew et al. (1996)
Meta-GGAs Perdew et al. (1999); Tao et al. (2003); Sun et al. (2015) introduce an extra dependence beyond the electron density by employing:
@@
where $\tau=1/2\sum_{i}^{\mathrm{occ}}|\nabla\varphi_{i}|^{2}$ is briefly referred to as the (KS) kinetic energy density. Kaplan and Perdew (2022) Here we focus on those MGGAs that depend on $\tau$ via:
@@
where $\tau_{\rm unif}=\frac{3}{10}(3\pi^{2})^{2/3}\,n^{5/3}$ is the kinetic energy density of the noninteracting Fermi gas at density $n$. In particular, the SCAN form is defined by setting $F_{\mathrm{xc}}^{\mathrm{SCAN}}(n,s,\alpha)=h_{\mathrm{xc}}^{1}(n,s)+\left[h_{\mathrm{xc}}^{0}(n,s)-h_{\mathrm{xc}}^{1}(n,s)\right]f_{\mathrm{xc}}(\alpha)$ where $f_{\mathrm{xc}}(\alpha)$ interpolates between the semi-local xc enhancement factors for single-orbital densities ($h_{\mathrm{xc}}^{0}$) and for slowly-varying densities ($h_{\mathrm{xc}}^{1}$). This modeling exploits the fact that $\alpha$ [see Eq. (6) ] tends to zero in the iso-orbital limit (i.e., when the density is dominated by a single occupied orbital), tends to one in the uniform-density limit, and becomes $\gg 1$ in regions of low density between closed shells.
From DFT to Spin-DFT.  For solving ground states involving collinear spin polarization either induced by external Zeeman couplings or spontaneously via magnetic phase transitions, the standard approach is to switch from DFT to Spin-DFT.von Barth and Hedin (1972); Gunnarsson and Lundqvist (1976) The method requires to solve a spin-dependent, single-orbital problem:
@@
Extended forms for exchange can be readily generated directly from the DFT approximations using the
$E_{\rm x}^{\rm DFA}[n_{\uparrow},n_{\downarrow}]=\frac{1}{2}E_{\rm x}^{\rm DFA}[2n_{\uparrow}]+\frac{1}{2}E_{\rm x}^{\rm DFA}[2n_{\downarrow}]\;.$ (8)
For correlation we must use:
$E_{{\rm c}}^{\rm DFA}[n_{\uparrow},n_{\downarrow}]=\int d^{3}r~{}n~{}\varepsilon^{\rm unif}_{\rm c}(n,\zeta)~{}F^{\rm DFA}_{\rm c}(n,\zeta,s,\alpha)\;;$ (9)
more in detail, unlike exchange, the correlation functional depends in a non separable fashion on $n_{\uparrow}$ and $n_{\downarrow}$ through the spin polarization (density) $\zeta=(n_{\uparrow}-n_{\downarrow})/(n_{\uparrow}+n_{\downarrow})$. Also note that the exchange functional of Eq. (8), $E_{\rm x}^{\rm DFA}[n_{\uparrow},n_{\downarrow}]$, can be equivalently reexpressed as $E_{\rm x}^{\rm DFA}[n,\zeta]$ in terms of the total density $n$ and polarization $\zeta$.
Issues.  Notice that the crucial quantity $\alpha$ does not carry any explicit information on $\nabla\zeta$. Therefore, the above expressions for GGAs and MGGAs come with a striking asymmetry. While the gradients of $\zeta$ are included in the exchange functional indirectly via the spin-scaling relation [i.e., Eq. (8)], the correlation form has no dependence on $\nabla\zeta$.
Considering how to include the physical dependence of the functional on $\nabla\zeta$ made us realize that a more fundamental asymmetry is embedded in this approach: neither the GGAs nor the MGGAs are invariant under U(1)$\times$×SU(2) gauge transformations. Although this deficiency might not be perceived as a problem as long as we consider scenarios that are not gauge symmetric — such as ferromagnets or antiferromagnets — one should not confuse the symmetry properties of particular solutions with those of the exact functional. To sort the latter out, we must enlarge our modeling framework.
From Collinear Spin-DFT to SCDFT.  Therefore, let us account not only for a scalar multiplicative potential $v_{\rm ext}$ and a multiplicative magnetic field ${B}^{a}$, but also for a (particle-) vector potential ${\bf A}$ and a spin-vector potential ${\bf A}^{a}$not (b) While ${\bf A}$ is useful to represent the vector potential of an external magnetic field, ${\bf A}^{a}$ is useful to represent the (one-body) spin-orbit couplings in the system. Vignale and Rasolt (1988); Bencheikh (2003)
The KS equations in SCDFT have the form of single-particle Pauli equations: Vignale and Rasolt (1988); Bencheikh (2003)
@@
Because $E_{\rm xc}[n,{\vec{m}},{\bf j},{\vec{\bf J}}]$ must yield the exact interacting densities via the KS densities, it must also be form invariant under (local) U(1)$\times$×SU(2) gauge transformations.Vignale and Rasolt (1988); Bencheikh (2003); Pittalis et al. (2017) In detail, such a transformation acts on spinors as follows:
@@
Revisiting the SCAN Form.  We recall from Eq. (5) that MGGAs, such as SCAN, explicitly depend on the KS kinetic energy density $\tau$. For dealing with non-collinear (nc) solutions, the kinetic energy density $\tau$ must be expressed in terms of the occupied two-component KS spinors $\tau_{\rm nc}=1/2\sum_{k}^{\mathrm{occ}}\Big{(}\partial_{\mu}\Phi^{\dagger}_{k}\Big{)}\Big{(}\partial_{\mu}\Phi_{k}\Big{)}$. But, it can be directly verified that $\tau_{\rm nc}$ is not gauge invariant.Pittalis et al. (2017) Thus, $\alpha_{\rm nc}$ obtained by replacing $\tau$ with $\tau_{\rm nc}$ in Eq. (6) is not invariant either. The corresponding extension of the SCAN functional by letting $\alpha\to\alpha_{\rm nc}$ would therefore lack the most fundamental property for dealing with magnetism. Fortunately, this problem can be fixed by the replacement ${\tau}_{\rm nc}\rightarrow\widetilde{\tau}_{\rm nc}$ with
@@
where ${\tau}^{a}_{\rm nc}=1/2\sum_{k}^{\rm occ}\Big{(}\partial_{\mu}\Phi^{\dagger}_{k}\Big{)}{\sigma}^{a}\Big{(}\partial_{\mu}\Phi_{k}\Big{)}\;$ is the spin-kinetic energy density. Eq. (12) was derived — among a few possible others — in previous work on the structure of the exact-exchange hole (see Ref. Pittalis et al., 2017), which incidentally also motivated the inclusion of the nc (spin-)kinetic energy density as a variable on which nc semi-local approximations must depend. Eq. (12) not only includes $\nabla\vec{m}$ and the (spin-)paramagnetic currents but, for consistency with the exact gauge-invariance requirements, also the coupling between ${\vec{m}}$ and ${\vec{\tau}}_{\rm nc}$.
Notice that Eq. (12) includes the well-known particle current corrections that ensure the U(1) invariance of the MGGA.Tao and Perdew (2005) It also includes the spin current corrections that were recently shown to ensure the SU(2) invariance of the MGGA for time-reversal symmetric spin-orbit coupling.Desmarais et al. (2024a) Eq. (12), however, is more general than the previous expressions as it also ensures invariance under conjoint U(1)$\times$×SU(2) transformations.Desmarais et al. (2024b)
Restricting ourselves to states with vanishing currents, the collinear (coll) limit of Eq. (12) is given by
$\widetilde{\tau}_{\rm nc}\rightarrow\widetilde{\tau}_{\rm coll}=\tau+\frac{{m}^{z}{\tau}^{z}}{n}+\frac{\nabla{m}^{z}\cdot\nabla{m}^{z}}{8n}\;.$ (13)
In what follows we will make use of $n=n^{\uparrow}+n^{\downarrow}$, $m^{z}=n^{\uparrow}-n^{\downarrow}$, $\tau=\tau^{\uparrow}+\tau^{\downarrow}$ and $\tau^{z}=\tau^{\uparrow}-\tau^{\downarrow}$. Therefore, the following magnetic modification of the SCAN functional (namely, mSCAN) suggests itself:
$E^{\rm mSCAN}_{x}[n_{\uparrow},n_{\downarrow}]=\int d^{3}r~{}n~{}\epsilon^{\rm unif}_{x}\left(n_{\uparrow},n_{\downarrow}\right)~{}F^{\rm SCAN}_{x}\left(n,s,\widetilde{\alpha}\right)$ (14a)
@@
$E^{\rm mSCAN}_{c}[n_{\uparrow},n_{\downarrow}]=\int d^{3}r~{}n~{}\epsilon^{\rm unif}_{c}\left(n_{\uparrow},n_{\downarrow}\right)~{}F^{\rm SCAN}_{c}\left(n,\zeta,s,\widetilde{\alpha}\right)$ (14b)
where
@@
with $\tau^{\uparrow}_{\rm unif}=3/5(6\pi^{2})^{2/3}(n^{\uparrow})^{5/3}$. Eq. (15) defines a quantity that is positive definite and, for closed-shell systems, $\widetilde{\alpha}=\alpha$. Thus, in Eq. (14a) we use the closed-shell exchange-enhancement factor of SCAN and in Eq. (14b) the correlation-enhancement factor of SCAN. But they are evaluated on $\widetilde{\alpha}$ rather than on $\alpha$.
Applications. We implemented the mSCAN functional in a developer’s version of the Crystal23 program.Erba et al. (2023) We study current-less, collinear magnets. Thus, the single-particle equation of SCDFT given in Eq. (10) is considerably simplified, as the vector potentials ${\bf A}_{\rm xc}$ and ${\bf A}^{a}_{\rm xc}$ vanish, and we must self-consistently solve the simpler Eq. (7) in a basis of Bloch orbitals $\varphi_{i,\sigma,\mathbf{k}}$ (with wavevector $\mathbf{k}$) formed from a linear combination of atom-centered spherical Gaussian-type functions (GTF).Desmarais et al. (2018)
In Eq. (7) for mSCAN, in a generalized KS context, the spin-dependent xc potential $v_{{\rm xc},\sigma}$ reads:Neumann et al. (1996)
@@
We solve Eqs. (7) and (16). Computational details are provided in the Supplementary Material (see also references Vilela Oliveira et al., 2019; Peterson et al., 2007; Saunders et al., 1992; Towler et al., 1996; Doll et al., 2001, 2006; Desmarais et al., 2023 therein).ESI
Table 1: Spin magnetic moments (in $\mu_{b}$) and, within parentheses, magnetic energies (in meV) of ferromagnetic metals. Results are shown from DFAs belonging to the three rungs of semi-local approximations of the Jacob’s ladder, and compared with experimental data on spin magnetic moments.

### v2

@@ Abstract
The Jacob’s ladder of density functional theory (DFT) proposes the compelling view that by extending the form of successful approximations — being guided by exact conditions and selected (least empirical) norms — upper rungs will do better than the lower, thus allowing to balance accuracy and computational effort. Meta-generalized-gradient-approximations (MGGAs) belong to the last rung of the semi-local approximations before hybridization with non-local wave function theories. Among the MGGAs, the Strongly Constrained and Appropriately Normed Approximation (SCAN) greatly improves upon GGAs from the lower rung. But the over magnetized solutions of SCAN make GGAs more reliable for magnetism. Here, we provide a solution that satisfies the most pressing desiderata for a density functional approximations for ferromagnetic, antiferromagnetic and non-collinear states. The approach is available in an implementation in the Crystal electronic structure package.
@@
Introduction.  Despite nearly one hundred years of quantum physics, magnetism continues to challenge our understanding of electron correlations. An unsettling example is given by the recently developed strongly-constrained and appropriately normed (SCAN) meta-generalized gradient approximation (MGGA) for density functional theory (DFT).Sun et al. (2015, 2016)
@@
SCAN satisfies a large set of exact constraints, from known properties of the exact exchange-correlation (xc) functional and gives accurate results at a computational cost of about a generalized-gradient-approximation (GGA). Thus, it can be considered a modern milestone from which it is possible to develop more accurate approximations.Aschebrock and Kümmel (2019); Lebeda et al. (2024)
@@
But SCAN does not reproduce correctly the stability and properties of prototypical ferromagnets (FM) like Fe. Fu and Singh (2018); Ekholm et al. (2018); Fu and Singh (2019); Mejía-Rodríguez and Trickey (2019); Tran et al. (2020) Overestimated tendency toward magnetism and exaggeration of magnetic energies are obtained for transition metal systems. Extension of the analysis to antiferromagnetic (AFM) solids has shown that modifications of MGGAs yielding an improved description of FM metals produce a deterioration in AFM insulators, and vice versa.Tran et al. (2020) The status is such that MGGAs yield no advantage over lower rungs of the semi-local approximation.Romero and Verstraete (2018); Tran et al. (2020)
@@
Here, we argue that the problem has persisted because the most important exact constraint for dealing with magnetism has not guided most of the DFT developments thus far. We resolve this issue by switching from collinear spin-DFTvon Barth and Hedin (1972); Gunnarsson and Lundqvist (1976) (the current standard methodology) to the fully non-collinear (nc) Vignale-Rasolt-Bencheikh spin-current DFT (SCDFT) – a complete DFT theory for dealing with magnetism within the framework of the two-component Pauli equation.Vignale and Rasolt (1988); Bencheikh (2003) As will be shown, this extended framework can also solve problems for collinear DFT.
@@
In SCDFT, the most distinctive property of the exact xc energy functional is the form invariance under U(1)$\times$×SU(2) gauge transformations.Vignale and Rasolt (1988); Bencheikh (2003); Pittalis et al. (2017) These transformations rotate spinors as well as their phase at each point in space [see Eq. (11) below]. Failing to satisfy this invariance leads to functional forms that do not necessarily distinguish between physical variations of the spin density and those that arise from a mere change of reference frame. The gauge invariance of the exact xc-energy functional is thus the most important exact constraint for guiding the construction of magnetic approximations.
Working within the general SCDFT, extra building blocks for the construction of extended approximate functional forms can be gained.Pittalis et al. (2017) In this work, we shall see, the collinear isoorbital indicators used in the SCAN (more below), which treat spin-up and spin-down channels separately, yield unbalanced magnetic solutions, Mejía-Rodríguez and Trickey (2019); Tran et al. (2020) and can be replaced by a unified non-collinear quantity normed by the gauge principle. As a result, the major issues observed in the collinear SCAN solutions get resolved, and the revised functional form is seamlessly extended to the more general non-collinear regime.
@@
From DFT to collinear Spin-DFT. DFT can in principle find the particle density and total energy of real interacting ground states by solving (self-consistently) an auxiliary non-interacting problem
@@
which is the celebrated Kohn-Sham equation. This is a Schrödinger-like equation that describes single-particle orbitals corresponding to the actual external potential (the second term on the l.h.s.) added to the Hartree potential (the third term on the l.h.s.) and the xc potential, $v_{\rm xc}$. The latter potential is such that the actual particle density is retrieved $n=\sum^{occ}_{i}|\varphi_{i}|^{2}$not (a) Correspondingly, the energy of the ground state of the real system $E$ is thus obtained by
@@
where $T_{s}=\frac{1}{2}\sum_{i}^{\rm occ}\int d^{3}r~{}|\nabla\varphi_{i}|^{2}$ is the Kohn-Sham kinetic energy, $E_{\rm H}[n]=\int\int d^{3}rd^{3}r^{\prime}\frac{n({\bf r})n({\bf r}^{\prime})}{2|{\bf r}-{\bf r}^{\prime}|}$ is the Hartree energy and $E_{\rm xc}[n]$ is the xc energy functional. For retrieving the exact total energy of the actual state, $E_{\rm xc}[n]$ accounts for the electron-electron energy beyond the classical evaluation obtained via $E_{\rm H}[n]$ as well as the kinetic energies beyond $T_{s}$. Also notice that $v_{\rm xc}=\delta E_{\rm xc}/\delta n$. All that is needed to make the above approach practical is a density functional approximation (DFA) for $E_{\rm xc}[n]$. Solutions to $n$ and $E$ can thus be found in an iterative fashion.
The first and simplest $E^{\rm DFA}_{\rm xc}[n]$ is the local density approximation (LDA) Kohn and Sham (1965)
@@
where the xc energy density is approximated locally with that of the uniform gas, $\epsilon^{\rm unif}_{\rm xc}(n)$, for which analytical and accurate numerical reference data are available. To capture the variations of more realistic states, a GGA uses information retrieved from the gradients of the density via $s=|\nabla n|/\left[2(3\pi^{2})^{1/3}n^{4/3}\right]$
@@
where $F_{\rm xc}^{\rm GGA}(n,s)$ is the xc enhancement factor that connects the LDA energy densities to the GGA energy densities. For slowly varying density, the LDA is retrieved (i.e. $F_{\rm xc}^{\rm GGA}(n,0)=1$). The most popular choice for the enhancement factor is the one by Perdew, Burke and Ernzerhof (PBE) $F_{\rm xc}^{\rm PBE}$. Perdew et al. (1996)
MGGAs Perdew et al. (1999); Tao et al. (2003); Sun et al. (2015) introduce an extra dependence beyond the electron density by employing
@@
where $\tau=1/2\sum_{i}^{\mathrm{occ}}|\nabla\varphi_{i}|^{2}$ is the (KS) kinetic energy density. Kaplan and Perdew (2022) Here we focus on those MGGAs that depend on $\tau$ via
@@
where $\tau_{\rm unif}=\frac{3}{10}(3\pi^{2})^{2/3}\,n^{5/3}$ is the kinetic energy density of the uniform gas at density $n$. In particular, the SCAN form is defined by setting $F_{\mathrm{xc}}^{\mathrm{SCAN}}(n,s,\alpha)=h_{\mathrm{xc}}^{1}(n,s)+\left[h_{\mathrm{xc}}^{0}(n,s)-h_{\mathrm{xc}}^{1}(n,s)\right]f_{\mathrm{xc}}(\alpha)$ where $f_{\mathrm{xc}}(\alpha)$ interpolates between the semi-local xc enhancement factors for single-orbital densities ($h_{\mathrm{xc}}^{0}$) and for slowly-varying densities ($h_{\mathrm{xc}}^{1}$). This modeling exploits the fact that $\alpha$ [see Eq. (6) ] tends to zero in the iso-orbital limit (i.e., when the density is dominated by a single occupied orbital), tends to one in the uniform-density limit, and becomes $\gg 1$ in regions of low density between closed shells.
For solving ground states involving collinear spin polarization either induced by external Zeeman couplings or spontaneously via magnetic phase transitions, the standard approach is to switch from DFT to Spin-DFT.von Barth and Hedin (1972); Gunnarsson and Lundqvist (1976) The method requires to solve a spin-dependent, single-orbital problem
@@
Extended forms for exchange can be generated from given DFT approximations using the collinear spin decomposition for exchange not (b)
$E_{\rm x}^{\rm SDFA}[n_{\uparrow},n_{\downarrow}]=\frac{1}{2}E_{\rm x}^{\rm DFA}[2n_{\uparrow}]+\frac{1}{2}E_{\rm x}^{\rm DFA}[2n_{\downarrow}]\;.$ (8)
For correlation, we must use
$E_{{\rm c}}^{\rm SDFA}[n_{\uparrow},n_{\downarrow}]=\int d^{3}r~{}n~{}\varepsilon^{\rm unif}_{\rm c}(n,\zeta)~{}F^{\rm SDFA}_{\rm c}(n,\zeta,s,\alpha)\;.$ (9)
Unlike exchange, the correlation functional depends in a non separable fashion on $n_{\uparrow}$ and $n_{\downarrow}$ through the spin polarization (density) $\zeta=|(n_{\uparrow}-n_{\downarrow})|/(n_{\uparrow}+n_{\downarrow})$; $\alpha=(\tau-\frac{|\nabla n|^{2}}{8n})/\tau_{\rm unif}d_{s}$ is the spin-summed iso-orbital indicator with $\tau=1/2\sum_{i,\sigma}^{\mathrm{occ}}|\nabla\varphi_{i}|^{2}$ and $d_{s}=\frac{1}{2}[(1+\zeta)^{5/3}+(1-\zeta)^{5/3}]$. Also note that the exchange functional of Eq. (8), $E_{\rm x}^{\rm DFA}[n_{\uparrow},n_{\downarrow}]$, can be equivalently reexpressed as $E_{\rm x}^{\rm DFA}[n,\zeta]$ in terms of the total density $n$ and polarization $\zeta$.
Issues.  Notice that $\alpha$ does not carry any explicit information on $\nabla\zeta$. Therefore, the above expressions for GGAs and MGGAs come with a striking asymmetry. While the gradients of $\zeta$ are included in the exchange functional indirectly via the spin-scaling relation [i.e., Eq. (8)], the correlation form has no dependence on $\nabla\zeta$.
This asymmetry reflects the lack of invariance of the approximate forms under U(1)$\times$×SU(2) gauge transformations. Although this deficiency might not be perceived as a problem as long as we consider scenarios that are not gauge symmetric — such as FMs or AFMs — one should not confuse the symmetry properties of particular solutions with those of the exact functional. To sort the latter out, we must enlarge our modeling framework.
Switching to SCDFT.  Let us account not only for a scalar multiplicative potential $v_{\rm ext}$ and a multiplicative magnetic field ${B}^{a}$, but also for a (particle-) vector potential ${\bf A}$ and a spin-vector potential ${\bf A}^{a}$not (c) While ${\bf A}$ is useful to represent the vector potential of an external magnetic field, ${\bf A}^{a}$ is useful to represent the (one-body) spin-orbit couplings in the system. Vignale and Rasolt (1988); Bencheikh (2003)
The KS equations in SCDFT have the form of single-particle Pauli equations Vignale and Rasolt (1988); Bencheikh (2003)
@@
Because $E_{\rm xc}[n,{\vec{m}},{\bf j},{\vec{\bf J}}]$ must yield the exact interacting densities via the KS densities, it must also be form invariant under (local) U(1)$\times$×SU(2) gauge transformations.Vignale and Rasolt (1988); Bencheikh (2003); Pittalis et al. (2017) In detail, such a transformation acts on spinors as follows
@@
We recall from Eq. (5) that MGGAs, such as SCAN, explicitly depend on the KS kinetic energy density $\tau$. For dealing with nc solutions, the kinetic energy density $\tau$ must be expressed in terms of the occupied two-component KS spinors $\tau_{\rm nc}=1/2\sum_{k}^{\mathrm{occ}}\Big{(}\partial_{\mu}\Phi^{\dagger}_{k}\Big{)}\Big{(}\partial_{\mu}\Phi_{k}\Big{)}$. But, it can be directly verified that $\tau_{\rm nc}$ is not gauge invariant.Pittalis et al. (2017) Thus, $\alpha_{\rm nc}$ obtained by replacing $\tau$ with $\tau_{\rm nc}$ in Eq. (6) is not invariant either. The corresponding extension of the SCAN functional by letting $\alpha\to\alpha_{\rm nc}$ would therefore lack the most fundamental property for dealing with magnetism. Fortunately, this problem can be fixed by the replacement ${\tau}_{\rm nc}\rightarrow\widetilde{\tau}_{\rm nc}$ withPittalis et al. (2017); Desmarais et al. (2024a)
@@
where ${\tau}^{a}_{\rm nc}=1/2\sum_{k}^{\rm occ}\Big{(}\partial_{\mu}\Phi^{\dagger}_{k}\Big{)}{\sigma}^{a}\Big{(}\partial_{\mu}\Phi_{k}\Big{)}\;$ is the spin-kinetic energy density.
Eq. (12) not only includes $\nabla\vec{m}$ and the (spin-)paramagnetic currents but, for consistency with the exact gauge-invariance requirements, also the coupling between ${\vec{m}}$ and ${\vec{\tau}}_{\rm nc}$.
Eq. (12) includes the well-known particle current corrections that ensure the U(1) invariance of the MGGA, which has been advanced both in practice and in theory.Dobson (1993); Becke (2002); Tao and Perdew (2005); Pittalis et al. (2006, 2007); Räsänen et al. (2009); Furness et al. (2015); Tellgren et al. (2018); Laestadius et al. (2019); Pemberton et al. (2022) It also includes the spin current corrections that were recently shown to ensure the SU(2) invariance of the MGGA for time-reversal symmetric spin-orbit coupling — the JSCAN.Desmarais et al. (2024b) Eq. (12) was also at the center of the derivation of a non-collinear ELF.Desmarais et al. (2024a) But it did not find exploitation in energy functional approximations to determine magnetic solutions.
We recall that the problems of overmagnetization by SCAN were imputed, in Refs. Mejía-Rodríguez and Trickey, 2019; Tran et al., 2020, to the different behaviour of the isoorbital indicators in the exchange of SCAN for the two spin channels. Below, we solve these problems by invoking a single $\widetilde{\alpha}$ for both spin channels — normed by the gauge principle. For capturing energies rather than just bond pictures, the following gauge invariant extension of SCAN suggests itself
$E^{\rm nc}_{x}[n,\vec{m},\mathbf{j},\vec{\mathbf{J}}]=\int d^{3}r~{}n~{}\varepsilon^{\rm unif}_{x}\left(n,\zeta_{\rm nc}\right)~{}F^{\rm SCAN}_{x}\left(n,s,\widetilde{\alpha}_{\rm nc}\right)\;,$ (13a)
@@
$E^{\rm nc}_{c}[n,\vec{m},\mathbf{j},\vec{\mathbf{J}}]=\int d^{3}r~{}n~{}\varepsilon^{\rm unif}_{\rm c}(n,\zeta_{\rm nc})~{}F^{\rm SCAN}_{\rm c}(n,\zeta_{\rm nc},s,\widetilde{\alpha}_{\rm nc})\;,$ (13b)
where $\zeta_{\rm nc}\equiv|\vec{m}|/n$, and
$\widetilde{\alpha}_{\rm nc}\equiv\frac{2n\widetilde{\tau}_{\rm nc}-\frac{(\nabla n)\cdot(\nabla n)}{4}}{\left(n+|\vec{m}|\right)\tau_{\rm unif}^{+}+\left(n-|\vec{m}|\right)\tau_{\rm unif}^{-}}$ (14)
in which $\tau_{\rm unif}^{\pm}=3/10(3\pi^{2})^{2/3}(n\pm|\vec{m}|)^{5/3}$, and all uniform gas quantities have been upgraded to an nc form by locally quantizing spin along the axis defined by the spin density $\vec{m}$.Kübler et al. (1988); not (d) The derivation of Eqs. (14) and (15) is provided in the Supplementary Material (this includes Ref.s Giuliani and Vignale, 2005; Lehtola et al., 2020; Desmarais et al., 2021; Vilela Oliveira et al., 2019; Peterson et al., 2007; Dolg, 2005; Balabanov and Peterson, 2005; Wilson et al., 1996; Campetella et al., 2020; Saunders et al., 1992; Towler et al., 1996; Doll et al., 2001, 2006; Desmarais et al., 2023; Yang et al., 2018; Wenzel et al., 2022; El Idrissi et al., 1991; Jones et al., 2022; Huber, 1979).ESI ; not (e)
For self-consistent collinear solutions, a modified SCAN functional (named mSCAN) is straightforwardly obtained solely in terms of $n^{\sigma}$, $\varepsilon^{\rm unif}_{\rm x/c}(n^{\uparrow},n^{\downarrow})$, and
@@
where $\tau^{\sigma}_{\rm unif}=3/5(6\pi^{2})^{2/3}(n^{\sigma})^{5/3}$.not (f) Finally, the mSCAN reduces to the SCAN on closed-shell solutions (more below).
Notable magnetic SCAN failures — resolved. Let’s scrutinize the novel collinear solutions. We implemented the mSCAN functional in a developer’s version of the Crystal23 program.Erba et al. (2023) We study current-less, collinear magnets. Thus, the single-particle equation of SCDFT given in Eq. (10) is considerably simplified, as the vector potentials ${\bf A}_{\rm xc}$ and ${\bf A}^{a}_{\rm xc}$ vanish, and we must self-consistently solve the simpler Eq. (7) in a basis of Bloch orbitals $\varphi_{i,\sigma,\mathbf{k}}$ (with wavevector $\mathbf{k}$) formed from a linear combination of atom-centered spherical Gaussian-type functions (GTF).Desmarais et al. (2018) In Eq. (7) for mSCAN, in a generalized KS context,Seidl et al. (1996); Desmarais et al. (2024c) the spin-dependent xc potential $v_{{\rm xc},\sigma}$ readsNeumann et al. (1996)
@@
We solve Eqs. (7) and (16). Computational details are provided in the SM.
Before applying mSCAN to prototypical magnets, notice that the x-mSCAN does not fulfill the (strictly collinear) spin decomposition relation of Eq. (8), yet it fulfills the exact linear scaling conditions (collinear or non-collinear). An extended analysis is provided in the SM.
[…]
Table 1: Spin moments (in $\mu_{b}$/atom) and, within parentheses, magnetic energies (in meV) of FM metals as calculated with DFAs and compared against experiments.